I am three times as old as my son. Five years later I shall be two and a half times as old as my son. What is my age?
Answer by JulietG(1812)

F = 3S [I am three times as old as my son.]
F + 5 = 2.5(S+5) [Five years later I shall be two and a half times as old as my son.]
Substitute the known value of F from the first equation into the second.
(3S) + 5 = 2.5(S+5)
Distribute
3S + 5 = 2.5S + 12.5
Subtract 2.5S from each side
.5S + 5 = 12.5
Subtract 5 from each side
.5S = 7.5
Multiply each side by 2 in order to get a full S
S = 15
.
If the son is 15, the father is 45.
Let's plug it into the equation to make certain it's correct.
45 + 5 = 2.5 * 20
50 = 50
Success!